solve d2u/dt2 =9d2u/dx2 with 0《x《1

u(0,t)=0 , u(1,t)=0 , d u(x,0)/dt =0


u(x,0) = x 0《x《0.25

0.25 0.25《x《0.75

1-x 0.75《x《1


1
Expert's answer
2021-03-25T07:56:54-0400

The wave equation:

"\frac{d^2u}{dt^2}=c^2\frac{d^2u}{dx^2}"

The general solution:

"u(x,t)=\sum sin(n\pi x/l)(A_ncos(cn\pi t/l)+B_nsin(cn\pi t/l))"

"f(x)=u(x,0)=\sum A_nsin(n\pi x/l)"

"g(x)=u_t(x,0)=\sum B_n(cn\pi/l)sin(n\pi x/l)"

"A_n=\frac{2}{l}\displaystyle\intop ^l_0 f(x)sin(n\pi x/l)dx"

"B_n=\frac{2}{cn\pi}\displaystyle\intop ^l_0 g(x)sin(n\pi x/l)dx"

Then:

"B_n=0"

"A_n=2(\displaystyle\intop ^{0.25}_0 xsin(n\pi x)dx+0.25\displaystyle\intop ^{0.75}_{0.25} sin(n\pi x)dx+\displaystyle\intop ^{1}_{0.75} (1-x)sin(n\pi x)dx)"

"\displaystyle\intop ^{0.25}_0 xsin(n\pi x)dx=\frac{sin(n\pi x)-n\pi xcos(n\pi x)}{n^2\pi^2}|^{0.25}_0="

"=\frac{sin(n\pi /4)-0.25n\pi cos(n\pi /4)}{n^2\pi^2}"

"\displaystyle\intop ^{0.75}_{0.25} sin(n\pi x)dx=-\frac{cos(n\pi x)}{n\pi}|^{0.75}_{0.25}=-\frac{cos(3n\pi/4)-cos(n\pi/4)}{n\pi}"

"\displaystyle\intop ^{1}_{0.75} (1-x)sin(n\pi x)dx=(-\frac{cos(n\pi x)}{n\pi}-\frac{sin(n\pi x)-n\pi xcos(n\pi x)}{n^2\pi^2})|^{1}_{0.75}="

"=-\frac{cos(n\pi)-cos(3n\pi/4)}{n\pi}-\frac{-n\pi cos(n\pi )-sin(3n\pi /4)+0.75n\pi cos(3n\pi/4)}{n^2\pi^2}"


"u(x,t)=\sum A_nsin(n\pi x)cos(3n\pi t)"